OpenStudy (anonymous):

Find the point at which the line intersects the given plane. x = 1 + 4t, y = 4t, z = 2−3t ; x + 2y − z + 1 = 0 (x, y, z) = _____________?

OpenStudy (turingtest):

substitute the parameterization of each component, x, y, and z, into the equation of the plane and solve for t. then plug that t back into the 3 equations to determine the values of x, y, and z

OpenStudy (anonymous):

x = 1 + 4t y = 4t z = 2 - 3t now sub these back into the other equation x + 2y - z + 1 = 0 (1 + 4t) + 2(4t) - (2 - 3t) + 1 = 0 1 + 4t + 8t - 2 + 3t + 1 = 0 --- combine like terms 15t = 0 t = 0 now sub 0 in for t in the other equations x = 1 + 4t x = 1 + 4(0) x = 1 + 0 x = 1 y = 4t y = 4(0) y = 0 z = 2 - 3t z = 2 - 3(0) z = 2 - 0 z = 2 check... x + 2y - z + 1 = 0 1 + 2(0) - 2 + 1 = 0 1 - 2 + 1 = 0 2 - 2 = 0 0 = 0 (correct) x = 1, y = 0, and z = 2 any questions ?

OpenStudy (anonymous):

thank you